# ClutterDexGrasp: A Sim-to-Real System for General Dexterous Grasping in Cluttered Scenes

### v1

@@ Abstract
Dexterous grasping in cluttered scenes presents significant challenges due to diverse object geometries, occlusions, and potential collisions. Existing methods primarily focus on single-object grasping or grasp-pose prediction without interaction, which are insufficient for complex, cluttered scenes. Recent vision-language-action models offer a potential solution but require extensive real-world demonstrations, making them costly and difficult to scale. To address these limitations, we revisit the sim-to-real transfer pipeline and develop key techniques that enable zero-shot deployment in reality while maintaining robust generalization. We propose ClutterDexGrasp, a two-stage teacher-student framework for closed-loop target-oriented dexterous grasping in cluttered scenes. The framework features a teacher policy trained in simulation using clutter density curriculum learning, incorporating both a novel geometry- and spatially-embedded scene representation and a comprehensive safety curriculum, enabling general, dynamic, and safe grasping behaviors. Through imitation learning, we distill the teacher’s knowledge into a student 3D diffusion policy (DP3) that operates on partial point cloud observations. To the best of our knowledge, this represents the first zero-shot sim-to-real closed-loop system for target-oriented dexterous grasping in cluttered scenes, demonstrating robust performance across diverse objects and layouts. More details and videos are available at https://clutterdexgrasp.github.io/.
@@ 1 Introduction
To enable dexterous grasping in cluttered scenes, some works first generate a non-collision grasp pose dataset specifically for cluttered scenes, then train a model to generate grasp poses and execute via motion planning [15, 14, 13]. However, the grasp poses generated by such methods are open-loop and cannot handle unexpected changes. To achieve a closed-loop grasping policy, current methods rely on real-world imitation learning (IL)[16, 17, 18] or sim-to-real reinforcement learning (RL) [19, 20, 21]. However, real-world IL requires a large amount of human-collected data to generalize across diverse cluttered scenes[16, 22, 23], which is costly and lacks scalability. Current RL-based methods, which can achieve high success rates in the real world, are mainly designed for relatively simple scenes, such as single-object tabletop grasping [11, 12]. These methods, however, aren’t directly applicable to cluttered scenes because they rely on simplified object states, like 6D poses, to ensure stable and convergent teacher policy learning, yet fail to capture the essential local geometric details crucial for dexterous grasping in cluttered scenes [24, 25].
In this paper, we aim to develop a generalized closed-loop policy for dexterous target-oriented grasping in cluttered scenes, trained in a simulation environment without expert demonstrations [26, 19], and to achieve zero-shot sim-to-real transfer [27, 28, 29]. The main challenges are as follows: (1) training instability in RL due to high-dimensional observation and action spaces caused by the diversity of object geometries and layouts, as well as the high DoF of the dexterous hand [19, 21, 20], (2) the sim-to-real gap in observation space, policy safety, and physical dynamics [30, 31, 32]. To address these challenges, we propose ClutterDexGrasp, a two-stage teacher-student framework. The teacher is trained in simulation using clutter density curriculum learning to reduce the difficulty of directly learning complex grasping strategies. Additionally, we introduce a novel geometry and spatially embedded scene representation, which captures the local geometry of the target object and its surrounding environment, enabling the teacher policy to effectively learn human-like grasping strategies in cluttered scenes [1]. The student policy is trained using offline imitation learning with a 3D diffusion policy (DP3) [33] on partial point cloud observations, enabling zero-shot sim-to-real transfer. To ensure safety during contact-rich interactions, we introduce an interaction-aware safety curriculum that minimizes collisions and avoids harmful behaviors, such as excessive force.
@@ 2 Related Works
Dexterous grasping has progressed from single-object setups to complex cluttered environments [4, 5, 6, 3, 2]. Open-loop methods for cluttered scenes focus on grasp pose generation and planning [15, 14, 13, 4, 5], but lack adaptability. For dynamic control, closed-loop approaches using RL [7, 8, 9, 10, 11, 12, 35, 19, 21, 20, 36] achieve success in single-object scenes but face challenges in cluttered environments due to complex interactions and state representations [24, 25, 13, 14]. Alternately, IL-based methods [16, 17, 18, 37] require extensive demonstrations that are costly to collect for diverse cluttered scenarios [22, 23]. Recent advances using vision-language-action frameworks [16, 38] and motion capture [39] aim to reduce demonstration burden but still require significant human effort. Our approach addresses these challenges through RL with specialized representations and curriculum learning that can generalize to cluttered environments without real-world demonstrations.
@@ 4.1.1 Geometry and Spatial Representation
Novel Representation
Humans naturally grasp the target object in cluttered scenes with both grasp stability and collision avoidance. Inspired by this, we propose a novel representation that enables similar adaptive behavior. Our approach samples n points from the target object and m points from surrounding objects using their mesh files. For each finger joint, we compute distances to the nearest sampled points from both the target ($d_{pos}$) and non-target objects ($d_{neg}$), as illustrated in Fig. 9. This representation efficiently encodes both object geometry and spatial relationships between the hand and the clutter. It enables collision-aware grasping while maintaining a compact observation space that simplifies learning of complex geometric features. By incorporating this representation into both the observation space and reward function, we promote clutter-aware grasping without requiring computationally expensive point-cloud rendering. Detailed implementation is provided in Appendix B.1.

### v2

@@ Abstract
Dexterous grasping in cluttered scenes presents significant challenges due to diverse object geometries, occlusions, and potential collisions. Existing methods primarily focus on single-object grasping or grasp-pose prediction without interaction, which are insufficient for complex, cluttered scenes. Recent vision-language-action models offer a potential solution but require extensive real-world demonstrations, making them costly and difficult to scale. To address these limitations, we revisit the sim-to-real transfer pipeline and develop key techniques that enable zero-shot deployment in reality while maintaining robust generalization. We propose ClutterDexGrasp, a two-stage teacher-student framework for closed-loop target-oriented dexterous grasping in cluttered scenes. The framework features a teacher policy trained in simulation using clutter density curriculum learning, incorporating both a geometry- and spatially-embedded scene representation and a novel comprehensive safety curriculum, enabling general, dynamic, and safe grasping behaviors. Through imitation learning, we distill the teacher’s knowledge into a student 3D diffusion policy (DP3) that operates on partial point cloud observations. To the best of our knowledge, this represents the first zero-shot sim-to-real closed-loop system for target-oriented dexterous grasping in cluttered scenes, demonstrating robust performance across diverse objects and layouts. More details and videos are available at https://clutterdexgrasp.github.io/.
@@ 1 Introduction
To enable dexterous grasping in cluttered scenes, some works first generate a non-collision grasp pose dataset specifically for cluttered scenes, then train a model to generate grasp poses and execute via motion planning [15, 14, 13]. However, the grasp poses generated by such methods are open-loop and cannot handle unexpected changes. To achieve a closed-loop grasping policy, current methods rely on real-world imitation learning (IL)[16, 17, 18] or sim-to-real reinforcement learning (RL) [19, 20, 21]. However, real-world IL requires a large amount of human-collected data to generalize across diverse cluttered scenes[16, 22, 23], which is costly and lacks scalability. Current RL-based methods, which can achieve high success rates in the real world, are mainly designed for relatively simple scenes, such as single-object tabletop grasping [11, 12]. These methods, however, aren’t directly applicable to cluttered scenes because they rely on simplified object states, like 6D poses, to ensure stable and convergent teacher policy learning, yet fail to capture the essential local geometric details crucial for dexterous grasping in cluttered scenes [24, 25]. To compactly represent the interaction between the object and the dexterous hand, [26] proposes a representation that computes the distance between the links of the dexterous hand and the graspable and ungraspable areas, which further improves teacher policy training [27].
In this paper, we aim to develop a generalized closed-loop policy for dexterous target-oriented grasping in cluttered scenes, trained in a simulation environment without expert demonstrations [28, 19], and to achieve zero-shot sim-to-real transfer [29, 30, 31]. The main challenges are as follows: (1) training instability in RL due to high-dimensional observation and action spaces caused by the diversity of object geometries and layouts, as well as the high DoF of the dexterous hand [19, 21, 20], (2) the sim-to-real gap in observation space, policy safety, and physical dynamics [32, 33, 34]. To address these challenges, we propose ClutterDexGrasp, a two-stage teacher-student framework. The teacher is trained in simulation using clutter density curriculum learning to reduce the difficulty of directly learning complex grasping strategies. Additionally, we leverage the distance representation [26] and extend it to clutterd scenes by computing distance between links of the dexterous hand and the target and non-target objects, such a geometry and spatially embedded scene representation captures the local geometry of the target object and its surrounding environment, enabling the teacher policy to effectively learn human-like grasping strategies in cluttered scenes [1]. The student policy is trained using offline imitation learning with a 3D diffusion policy (DP3) [35] on partial point cloud observations, enabling zero-shot sim-to-real transfer. To ensure safety during contact-rich interactions, we introduce an interaction-aware safety curriculum that minimizes collisions and avoids harmful behaviors, such as excessive force.
@@ 2 Related Works
Dexterous grasping has progressed from single-object setups to complex cluttered environments [4, 5, 6, 3, 2]. Open-loop methods for cluttered scenes focus on grasp pose generation and planning [15, 14, 13, 4, 5], but lack adaptability. For dynamic control, closed-loop approaches using RL [7, 8, 9, 10, 11, 12, 37, 19, 21, 20, 38] achieve success in single-object scenes but face challenges in cluttered environments due to complex interactions and state representations [24, 25, 13, 14]. [26] and [27] introduce distance between links of the dexterous hand and graspable and ungraspable area of the target object as representation for RL training, with [27] demonstrating great performance on real robots using a teacher-student framework. However, their training is conducted exclusively on single objects, which limits their applicability to complex scenarios. In contrast, our approach directly incorporates cluttered scene representations during RL training, enabling enhanced performance in complex cluttered configurations, particularly under extreme stacking conditions. Alternately, IL-based methods [16, 17, 18, 39] require extensive demonstrations that are costly to collect for diverse cluttered scenarios [22, 23]. Recent advances using vision-language-action frameworks [16, 40] and motion capture [41] aim to reduce demonstration burden but still require significant human effort. Our approach addresses these challenges through RL with specialized representations and curriculum learning that can generalize to cluttered environments without real-world demonstrations.
@@ 4.1.1 Geometry and Spatial Representation
Cluttered Scene Representation
For each finger joint, we compute distances to the nearest sampled points from both the target ($d_{pos}$) and non-target objects ($d_{neg}$), as illustrated in Fig. 9. This representation efficiently encodes both object geometry and spatial relationships between the hand and the clutter. It enables collision-aware grasping while maintaining a compact observation space that simplifies learning of complex geometric features. By incorporating this representation into both the observation space and reward function, we promote clutter-aware grasping without requiring computationally expensive point-cloud rendering. Detailed implementation is provided in Appendix B.1.